OpenStudy (anonymous):

3^4 is 81 ? not sure :(

OpenStudy (bahrom7893):

3^4 = 3^(2+2) = 3^2 * 3^2 = 9*9 = 81

OpenStudy (tkhunny):

\(3^{4}\) MEANS 3*3*3*3. About what are you not sure?